A bicyclist rode into the country for 5 h. In returning, her
speed was 5 mi/h faster and the trip took 4 h. What was her speed each way?

Answer by checkley71(8403)

R5=(R+5)4
5R=4R+20
5R-4R=20
R=20 SPEED OF THE BIKE ON THE RIDE INTO THE COUNTRY.
PROOF
5*20=25*4
100=100